OpenStudy (anonymous):

Joey has a beach ball with a radius of 20 cm. What volume of air will it hold when fully inflated? [Round to nearest tenth.]

OpenStudy (anonymous):

\[20*20*20\]

OpenStudy (anonymous):

*pi

OpenStudy (anonymous):

or *3.14

OpenStudy (anonymous):

um what

OpenStudy (anonymous):

V=4/3pir^3 V=4/3*3.14*20^3 V=4.1867*8,000 V=33,493 cm^3 ans.